2. A particle moves so that its position vector at time t is given by ˜ r = e −t cost ˜ i + e −t sin t ˜ j. Show that at any time t, (a) its velocity ˜ v is inclined to the vector ˜ r at a constant angle 3π/4 radians. (b) its acceleration vector is at right angles to the vector ˜ r. 


1
Expert's answer
2021-04-22T10:59:29-0400

To be given in question

"\overrightarrow{r}" ="(e^{-t}cost) i+(e^{-t}sin) j"

To be asked in question

Velocity=?

Phase ="\frac{3 \pi}{4}"

r.a=?

We know that

"r=(e^{-t}cost )i+(e^{-t}sint)j \rightarrow(1)"


Equation differenciate with respect to t

"v=\frac{dr}{dt} \rightarrow(2)"


Put value in eqution (1)and(2)

"v=\frac{d}{dt}((e^{-t}cost) i+(e^{-t}sint) j)"

"v=(-e^{-t}cost-e^{-t}sint)i-(e^{-t}sint+e^{-t}cost)j \rightarrow(3)"

"a=(2e^{-t}sint)i+(-2e^{-t}cost)j"


eqution (1) take mode

"\mid r\mid =" "\sqrt{(e^{-t}cost)^2+(e^{-t}sint)^2}"

  1. "r^2=e^{-2t}cos^2t+e^{-2t}sin^2t+2e^{-2t}costsint =2e^{-2t}"

"|v|=\sqrt{2e^{-2t}}=\sqrt2e^-t"

"cos(r.v)=\frac{r.v}{|r||v|}=\frac{-e^{-2t}}{e^-t\sqrt2e^{-t}}=-\frac{\sqrt2}{2}"

"angle(r.v)=\frac{3\pi}{4}"

(b)

"r.a=e^{-t}cost(2e^{-t}sint)+e^{-t}sint(-2e^{-t}cost)=0"